the propeller of a World War 2 fighter plane is 2.1 m in diameter.

33% P art (a) What is its angular velocity in radians per second if it spins at 1280 rev/min? Grade Summary Deductions 09 Potential 1009 sin0 cotan0asin0 atan0 acotan sinh0 coshOtanh0cotanh0 cos0 Attempts remaining 4 5 6 per attempt) detailed view END VO CLEAR BACKSPACE Degrees , Radians Submit Hint I give up! Feedback: 0% deduction per feedback s: 0% deduction per hint. Hints remaining 33% Part (b) What is the linear speed (in ms) of its tip at this angular velocity if the plane is stationary on the tarmac? 33% Part(c) What is the centripetal acceleration of the propeller tip under these conditions? Represent your answer in meters per seconds are
Answer #1

a)

1 rpm=0.10472 rad/s

angular velocity in radians per second

W=1280*0.10472 =134.04 radians/second

b)

radius of propellor

r=D/2=2.1/2 =1.05 m

Linear speed

V=RW=1.05*134.04

V=140.74 m/s

c)

centripetal acceleration

a=V2/r =140.742/1.05

a=18865.42 m/s2